1. Translate into an equation and solve. (Use n to represent the smaller number.)
The sum of two numbers is twenty-four. Six less than five times the smaller is equal to the larger

2. Find three consecutive even integers whose sum is negative thirty-six.

Answer by ewatrrr(24785)

 
Hi,
1. smaller number n and larger number is (24-n)
Question states***
5n-6 = 24-n
6n = 30,
n = 30/6 = 5 & the larger number is 19
and 25-6+=+19
2. x + (x+2) + (x+4) = -36
3x = -42 , x = -14
-14,-12,-10 are the integers.